OpenStudy (anonymous):

use synthetic division to find p(-3) for p(x)=x4 -2x3 -4x +4

OpenStudy (anonymous):

283-54-12+4=221

OpenStudy (anonymous):

Do i just plug in the -3?

OpenStudy (anonymous):

yeah found it. :) subtract 151 from both sides. and use synthetic division to divide rightside from (x+3). yep now its a factor. write it like this p(x) -151 = (x+3)(x3 -5x2+15x--49) and when x =-3 right side is equal to 0; have fun

OpenStudy (anx):

I've included the answer in the pic but ill explain incase you can read my sloppy handwriting. \[p(x) = x ^{3}-5x ^{2}+15x-49+151\div (x+3)\] The degree loses one and is replaced in decending order. instead of x^4 being first, its now x^3. I can explain more if you still need help.

OpenStudy (anonymous):

I still need help with this one!

OpenStudy (anonymous):

Hi Minimoo, could you possibly rewrite your question? I'm currently having some difficulty understanding the problem. :)